### v1

# A Deep Neural Network Approach to Compact Source Removal

Abstract
Context. Analyzing extended emission in photometric observations of star-forming regions requires maps free from compact foreground, embedded, and background sources, which can interfere with various techniques used to characterize the interstellar medium. Within the framework of the NEMESIS project, we apply machine learning techniques to improve our understanding of the star formation timescales, which involves the unbiased analysis of the extended emission in these regions.
Aims. We present a deep learning-based method for separating the signals of compact sources and extended emission in photometric observations made by the Herschel Space Observatory, facilitating the analysis of extended emission and improving the photometry of compact sources.
Methods. Central to our approach is a modified U-Net architecture with partial convolutional layers. This method enables effective source removal and background estimation across various flux densities, using a series of partial convolutional layers, batch normalization, and ReLU activation layers within blocks. Our training process utilized simulated sources injected into Herschel images, with controlled flux densities against known backgrounds. A dynamic, signal-to-noise ratio-based adaptive masking system was implemented to assess how prominently a compact source stands out from the surrounding background.
Results. The results demonstrate that our method can significantly improve the photometric accuracy in the presence of highly fluctuating backgrounds. Moreover, the approach can preserve all characteristics of the images, including the noise properties.
Conclusions. The presented approach allows users to analyze extended emission without the interference of disturbing point sources or perform more precise photometry of sources located in complex environments. We also provide a Python tool with tutorials and examples to help the community effectively utilize this method.
Key Words.:
ISM: structure - Techniques: image processing - Techniques: photometric - Methods: data analysis
1 Introduction
Stars form as the result of a long series of processes that take place within dense molecular clouds. The first step of the process is the gravitational collapse (Shu, 1977; Shu et al., 1987; Larson, 2003). The resulting young stellar objects (YSOs) form young stellar groups or clusters, often organized in a “hub-filament structure”. André et al. (2014) proposed a paradigm suggesting that star formation occurs primarily within filaments, implying a universal process extending up to galactic scales. Further research has identified even smaller structures within filaments, known as “fibers” (Hacar et al., 2013, 2018, 2024; Lee et al., 2014; Zavagno et al., 2023), which exhibit similar dynamical properties across both low- and high-mass star-forming regions.
The structure of the interstellar medium (ISM) shows a great variety across the Milky Way. It is shaped by many physical processes, including heating, cooling, gravity and magnetic fields. We already know that star formation efficiency and star formation rates also show significant differences in different environments. Federrath & Klessen (2012) discussed that observations of star-forming clouds reveal that the star formation rate (SFR) surface density spans four orders of magnitude and positively correlates with the gas surface density (Heiderman et al., 2010), indicating that denser gas supports a higher rate of star formation. This raises a key question: How is gas locally compressed in the ISM to form dense cores that eventually become gravitationally unstable, leading to star formation? Gas compression through shocks, driven by large-scale supersonic turbulence, may play a critical — if not the primary — role in establishing the initial conditions for star formation (see, e.g., reviews by Mac Low & Klessen, 2004; McKee & Ostriker, 2007). This information can be inferred from the analysis of the extended emission.
To analyze extended emissions, maps need to be free from contaminating sources. Even in Fourier space, compact sources (those with extensions similar to or smaller than the beam’s FWHM at a specific wavelength) add substantial power to the image’s power spectrum across a broad range of spatial frequencies, affecting the image especially at frequencies near the beam size. Depending on the source density and clustering strength, lower spatial frequencies may also be contaminated, albeit with lower power density spread over a wide bandwidth.
Image analysis techniques include power spectrum analysis that quantifies the spatial distribution of fluctuations in the ISM by measuring how intensity varies across different spatial scales. It is widely used to study turbulence in the ISM, with different scales indicating the presence of large molecular clouds, small clumps, or filamentary structures. Power spectra help identify whether structures follow a particular scaling law (see, e.g. Roy et al., 2015).
The probability density functions (PDFs) represent the distribution of intensities (such as column densities or densities) in a region of the ISM, showing the frequency of different densities. PDFs are useful for distinguishing different physical regimes within the ISM, like a log-normal PDF often indicates turbulence-dominated gas, while power-law tails can suggest gravitational collapse or strong density contrasts, such as those in star-forming regions (see, e.g. Kirk et al., 2024).
$\Delta$-variance (Stutzki et al., 1998), introduced as a wavelet-based measure for the statistical scaling of structures in astronomical maps, is a tool for quantifying spatial scaling and structure, particularly helpful in characterizing turbulence. It measures how variance changes with scale, often providing more sensitivity to large-scale structures than power spectrum analysis. $\Delta$-variance is particularly effective for analyzing hierarchical structures and fractal-like features in the ISM, helping to differentiate between turbulent and non-turbulent regions. One of its applications to the analysis of extended emission is presented by Elia et al. (2014).
These are the most common techniques, but several others are used in the different sub-fields of astrophysics, recently discussed and compared by Metha & Berger (2024). In any case, comparing image analysis techniques becomes challenging without source subtraction, as different methods vary in their sensitivity to discrete substructures. Techniques relying on sparsity information can be skewed by just a few point sources, while methods analyzing entire intensity maps are more affected by clustering. Effective analysis of extended emission requires a source subtraction technique that targets a specific range of spatial frequencies while preserving the noise properties of the image. There have been several approaches for source extraction in the presence of structured background and for the removal of sources. Men’shchikov (2021) developed the getsf code, Molinari et al. (2011) the CuTEx code, and Marton et al. (2014) the boloSource() code, all of them using different approaches, but based on traditional mathematical and physical models.
In this paper, we present a deep-learning tool, designed to remove point sources from the photometric observations of the PACS instrument (Poglitsch et al., 2010) on-board of the Herschel Space Observatory (Pilbratt et al., 2010). This work is part of our Novel Evolutionary Model for the Early Stages of stars with Intelligent Systems (NEMESIS [^1] https://nemesis.univie.ac.at) project, and a continuation and extension of the new Herschel/PACS catalog (Marton et al., 2024). The tool can be trained for the photometric data of other instruments on other spacecrafts, as well, thanks to the modified U-Net architecture with partial convolutional layers integrated within an inpainting framework, central to our approach. In Section 2, we describe the data used for the different steps of preprocessing, training and validation. In Section 3, we give a detailed description of the deep-learning algorithm, the architecture, the implementation and the photometric methods used for validation. In Section 4, we present the performance of the method, and in Section 5, we discuss the photometric accuracy and the image properties of the resulting data after applying our method to various examples.
2 Data
In this section, we provide a detailed explanation of the data used in this study and its key properties. We describe how we created simulations for the task, how the dataset was split into training, validation, and test sets, the preprocessing steps we applied, and how we constructed the masks for the compact sources.
This problem requires careful preparation to approach it as a supervised learning problem. The main idea is straightforward: to remove an object from an image, we need to understand what lies behind it. To create training data, we use original images where the true background is already known and sources are then injected into these backgrounds, simulating the presence of sources as seen in real observational data. This approach provides paired data: the input image, which includes the added sources on the background, and the ground truth target image, which is the original background without the sources. It allows us to construct a dataset where the machine learning model can learn to predict the clean background from the input image containing compact sources. In the subsections that follow, we explain these steps in detail.
2.1 Herschel Science Archive (HSA)
The Herschel Space Observatory (Pilbratt et al., 2010) was the fourth cornerstone mission in the European Space Agency (ESA) science program. More than 35,000 observations were made during its mission. The different data products created by the Standard Product Generator (SPG) pipelines are available through the Herschel Science Archive (HSA) and can be downloaded via the web interface or through the HIPE software, specifically developed for Herschel data manipulation.
The PACS photometer was a dual-band instrument composed of two filled silicon bolometer arrays: the Blue camera, consisting of $32\times 64$ pixels, and a Red Camera, consisting of $16\times 32$ pixels. The Blue camera provided two filters centered at $70\mu\mathrm{m}$ and $100\mu\mathrm{m}$, (commonly referred to as BS - blue, and BL - green bands), while the Red camera acquired observations at the nominal wavelength of $160\mu\mathrm{m}$ (R - red band).
An important parameter of the PACS observations that we need to introduce is the Observing Mode, as we used two types of data. The Observing Mode could either be the nominal Scan Map or Parallel Mode. In the latter, the observations were acquired together with the SPIRE instrument with a simultaneous five-band photometry of the same field of view. In both modes, the bolometer read-out frequency was 40 Hz, but due to data-rate limitations, an on-board averaging of 4 frames was performed, except for the blue and green bands in Parallel Mode, where the averaging was increased to 8 frames. For our simulations, we used both Scan Map and Parallel Observing Modes for all the three bands, resulting in a total of six different types of data.
The standard satellite scanning speed was 20′′s-1. The highest speed of 60′′s-1 was dedicated to large Galactic surveys and Parallel Mode maps.
For our analysis, we exclusively utilized Level 2.5 and Level 3 maps created with the JScanam algorithm (Graciá-Carpio et al., 2017). These maps are well-suited for our purposes, as the JScanam method ensures reliable recovery of both extended emission and point-like sources without relying on filtering or noise models.
2.2 Point Spread Function
The shape of the PACS Point Spread Function (PSF) is dominated by a tri-lobe pattern in all three bands. A detailed description can be found in the PACS photometer point spread function technical note (Lutz, 2015). For our simulations, to simulate real sources, we used the official Vesta PSFs, which are specific to the Observing Mode, band, and scan speed (Lutz, 2015). We transformed the PSF images to match the resolution of the observations, resulting in the six PSF images, as shown in Fig. 1.
For 60′′s-1 scan speeds, the shape of the PSF is more elongated. However, with different scan angles, the elongation occurs in different directions, which can be handled using circular masks, as discussed in Section 2.4.4.
Figure 1: The figure shows the six Vesta PSF images after they have been transformed to match the resolution of the observation image data. The first row represents the Scan Map mode, the second is the Parallel mode, and the three columns correspond to the blue, green and red bands, respectively. Note that all the images were plotted to have the same size for visualization purposes, but it can be seen that the resolution of the PSF images differs between the bands and Observing Modes.
2.3 Simulations
As mentioned earlier, with the two Observing Modes and the three bands, we have a total of six different types of data to process. However, since the three bands within the modes are more similar in appearance in the image cutouts, we decided to merge the data for each mode. This way, we only need to train two models, one for the Scan Map mode and one for the Parallel mode, and allow the network to learn the differences between the bands using masks designed specifically for the bands, as described in Section 2.4.4. The need to handle Scan Map and Parallel modes separately arises from their resolutional differences.
2.3.1 Data Generation Process
We covered a wide range of spectral flux densities to ensure that our model could handle a variety of extreme cases. Flux density levels, defined from $10$ mJy to $40,000$ mJy, were chosen using irregular steps. Denser steps were used at lower flux densities, as they typically exhibit greater variation against the background compared to higher values, which show less visual difference. The categories are shown in Table 1. All the simulations were generated with the following steps:
  1. 1.
    An image field was selected with relatively few real sources present.
  2. 2.
    We determined how many random sources could be injected into the image to maximize the number of sources while avoiding large empty spaces and excessive clustering of sources in smaller areas. The choice of the number, ranging between 25 and 210, was loosely proportional to the image size, but manually adjusted for each map based on visual inspection.
  3. 3.
    We defined the minimum and maximum flux density levels to be used, so that at the minimum level, the injected sources would start to become visible against the background, and at the maximum level, the sources would be realistically bright compared to the background. This varied based on the complexity and brightness of the background in the selected image field.
  4. 4.
    We created as many simulation images for a specific field as there are flux density levels (defined in Table 1) within the determined minimum and maximum values (inclusive). For instance, suppose we determine to use flux levels between 40 and 125 for an image field. If we then check the table for values within this range, we find 40, 60, 80, 100, and 125 (inclusive). Therefore, five simulations will be created.
  5. 5.
    Since the PSF image arrays were normalized to have a flux density of exactly 1 Jy, we scaled it to match the desired flux density level by multiplying the values accordingly.
  6. 6.
    We generated the specified number of random positions for the particular field and injected the scaled PSFs into these positions.
  7. 7.
    A corresponding .csv file was also created, which included the equatorial coordinates of right ascension (Ra) and declination (Dec) in the J2000.0 reference frame of the injected sources, with the flux density levels used in mJy.
Table 1: The defined flux density levels from 10 mJy to 40,000 mJy that were used to create simulations.
Flux Density (mJy)
10 20 40 60 80 100 125 150 175 200 250 300 350
400 500 600 700 800 900 1000 1250 1500 1750 2000 2500 3000
3500 4000 4500 5000 6000 7000 8000 9000 10000 11000 12000 13000 14000
15000 16000 18000 20000 22500 25000 27500 30000 32500 35000 37500 40000
An example is shown in Fig. 2, where sources with flux density of 2,000 mJy were injected into the Scan Map Blue field with obsid=1342191801.
Figure 2: An example where a Scan Map Blue field with obsid=1342191801 (on the left) was used as an input, and sources with a flux density of $2,000$ mJy were injected at random positions (on the right).
2.3.2 Dataset Partitioning
Several steps have been taken to create a balanced and well-representative dataset. First, we listed the observation identifiers (obsids) of potential fields large enough to contain at least $25$ injected sources, and had only a few real sources present. We then sorted these obsids into a few custom categories defined by us, based on a subjective assessment of their similarity in terms of complexity, appearance and structure. From these categories, we selected the most suitable fields. See the six examples in Fig. 3, representing the six fields used in the Scan Map Blue test set.
Figure 3: The six fields used in the Scan Map Blue test dataset, categorized by approximate complexity. The upper left is the least complex, and complexity increases from left to right and top to bottom, with the bottom right being the most complex.
Next, we followed the steps described in Section 2.3.1 to create the simulations for each selected obsid field. We then divided the list of obsids into training, validation and test sets, ensuring that each of the three datasets contained at least one field from each of the custom categories. The training, validation and test sets are disjoint, since we used different obsids in each set. Table 2 shows the distribution of sources and fields in the Scan Map and Parallel datasets, and additional details about the fields used can be found in Table 6, Table 7 and Table 8 in the Appendix. The reason why we did not use simulations from the Parallel Red dataset for training is because it has the same PSF as the Scanmap version, thus only test simulations were created.
Table 2: Total number of sources and fields for BS, BL, and R in the Scan Map and Parallel datasets.
Scan Map Parallel
Number of Sources Number of Fields Number of Sources Number of Fields
BS BL R BS BL R BS BL R BS BL R
Training 72,570 72,345 64,390 29 29 23 61,345 35,080 N/A 16 6 N/A
Validation 14,755 17,095 14,430 7 7 4 15,230 9,485 N/A 4 3 N/A
Test 18,095 16,085 15,490 6 6 5 18,070 12,640 15,170 6 4 6
2.4 Preprocessing Workflow
We wanted our network to work on specified input coordinates rather than the entire image field. Image fields vary significantly in spatial dimensions, number of sources, and other properties, making them difficult to use directly as inputs. Additionally, to ensure photometric accuracy, the central coordinates of the sources were essential. Also, to prevent the model from unintentionally modifying parts of the image other than the compact sources, it was essential to focus the network’s operations locally around the specified coordinates. For these reasons, we used the given coordinates to extract smaller image cutouts centered on the compact sources. Below, we discuss the steps taken to prepare these inputs.
2.4.1 Image Cutout Preparation
For training, it is more efficient to prepare the image cutouts in advance and save them as .npy files, a binary format used by NumPy (Harris et al., 2020) for storing arrays efficiently. This approach saves time during training, as generating the cutouts on-the-fly would be time consuming. Instead, we can simply load the pre-saved .npy files during training. However, in evaluation and production modes, the input consists of only the entire image maps and specified coordinates, as this is the simplest format for users to prepare. The image cutouts are then generated dynamically during the workflow.
We experimented with different cutout window sizes, starting with smaller sizes, such as $32\times 32$ pixels, and then increasing them. Our aim was to find a size large enough to capture sufficient background information for the model to effectively extract features and remove even the brightest sources as well. We found that $96\times 96$ pixel images offer a good balance between network complexity, and providing enough background context for the network to successfully remove the sources and estimate the underlying background.
2.4.2 Normalization
In our preprocessing workflow, we chose per-image normalization instead of global normalization. The pixel intensities in our images represent flux values, and the dynamic range of these fluxes can vary significantly between images due to the presence of objects with vastly different brightness levels.
Global normalization, which applies the same scaling factors across the entire dataset resulted in worse results and performance during training compared to the per-image version. Extremely bright regions dominate the normalization process, causing the average or dimmer point sources and background information to be poorly represented. Logarithmic scaling did not improve the performance, with the network’s learning process remaining stagnant.
By using per-image normalization, each image is scaled individually based on its own minimum and maximum intensities. This allows the model to focus on the specific characteristics of each input image without being distorted by extreme flux variations present in the dataset. This approach also improves the network’s ability to generalize across different fields with varying intensity profiles, ensuring better overall performance.
2.4.3 Data Augmentation
Since the input coordinates may not always be perfectly centered on the source, we applied random translations, with a maximum shift of 2 pixels in both vertical and horizontal directions, to mitigate this issue. Additionally, we applied horizontal and vertical flips, as well as random rotations, to augment the data and introduce sources with varying orientations. This augmentation strategy helps improve the model’s robustness by introducing it to a wider variety of source positions and orientations. We implemented these transformations using PyTorch’s torchvision library (Paszke et al., 2019).
2.4.4 Determining Mask Size
For our setup, it was important to determine the appropriate shapes and sizes of the masks, which distinguish the regions of compact sources to be modified (removed) from the valid background.
At scan speeds of 60′′s-1, the PSF shape becomes more elongated. Due to varying scan angles, this elongation occurs in different directions, which can be addressed by employing circular masks. We experimented with single circular masks of different sizes, but this approach is not ideal. Bright sources, compared to their background, require larger masks, while faint objects, barely distinguishable from the background, need smaller masks. If a smaller mask is used for brighter sources, a significant portion of the PSF visible against the background will remain untreated and therefore not be removed. On the other hand, using large masks for faint sources covers more than necessary, reducing the valid region available for the network to work with.
To address this, we implemented dynamically calculated mask sizes for each source. To quantify how prominently a compact source stands out from the surrounding background, we calculated the signal-to-noise ratio (SNR) based on the peak flux of the source region. The peak intensity is effective for assessing contrast against the background, particularly for fainter sources. An illustration of the SNR calculation is shown in Fig. 4. Given the different resolutions between the Scan Map and Parallel modes, the sizes of the source circle and surrounding annulus were adjusted accordingly. Specifically, for the Scan Map mode ($1.6\times 1.6$ arcseconds per pixel), we used radii of 8 pixels for the source circle and 10 and 22 pixels for the inner and outer annulus, respectively. For the Parallel mode ($3.2\times 3.2$ arcseconds per pixel), we used 4 pixels for the source circle and 8 and 18 pixels for the annulus. The exception is the Parallel Red mode, which shares the same PSF as the Scan Map mode, so we applied the Scan Map parameters for the SNR calculations in that case.
This annulus provides a good estimate of the background by computing both the mean $\left(\mu_{\text{bg}}\right)$ and standard deviation $\left(\sigma_{\text{bg}}\right)$ of the pixel intensities within it. The SNR is then calculated as:
$SNR=\frac{F_{\text{peak}}-\mu_{\text{bg}}}{\sigma_{\text{bg}}}.$ (1)
Figure 4: An illustration of the SNR calculation, with red and blue circles indicating the source region and the surrounding annulus, respectively. For the examples, we visualized the Scan Map Blue PSF on a logarithmic scale on the left, and the Parallel Blue PSF on a logarithmic scale on the right.
A higher SNR indicates that the source is more distinguishable from the background, meaning a larger portion of the PSF is visible and influences the surrounding region.
We calculated the SNR values for $2,000$ random sources for each type of data and analyzed the relationship between the SNR values and the appropriate mask size for each PSF separately. The results showed that using six distinct mask sizes, based on SNR thresholds, provides enough flexibility to handle the main differences between sources while achieving good performance. Instead of mapping the SNR values directly to mask sizes with interpolation, we decided to use discrete thresholds. This is because background conditions can vary significantly, making the mapping between SNR and the appropriate mask size imprecise. For example, two sources with the same flux density and similar background conditions may still have different SNR values if one of them lies near a brighter cloud or another bright object within the annulus region. These slight fluctuations in SNR make it more reliable to use thresholded categories rather than a direct mapping from SNR to mask size. This approach also helps us handle extreme cases more effectively.
We named the six masks, from largest to smallest, as: Full Mask, Large Mask, Moderate Mask, Small Mask, Focused Mask, and Minimal Mask. However, the appropriate mask sizes vary across different bands, thus the radii for the six masks were defined separately for each data type. The specific values for these masks are provided in Table 3. As an example, Fig. 5 shows selected cases from the Scan Map Green dataset.
Table 3: Mask sizes (in pixel radius) corresponding to different SNR ranges for the datasets.
Scan Map Parallel
SNR Range Mask Name Blue Green Red SNR Range Mask Name Blue Green
SNR $\geq$ 300 Full Mask 12 13 10 SNR $\geq$ 855 Full Mask 9 10
140 $\leq$ SNR $<$ 300 Large Mask 11 11 9 590 $\leq$ SNR $<$ 855 Large Mask 8 9
13 $\leq$ SNR $<$ 140 Moderate Mask 9 10 8 300 $\leq$ SNR $<$ 590 Moderate Mask 7 8
8 $\leq$ SNR $<$ 13 Small Mask 7 8 7 20 $\leq$ SNR $<$ 300 Small Mask 6 7
4 $\leq$ SNR $<$ 8 Focused Mask 6 6 6 13 $\leq$ SNR $<$ 20 Focused Mask 5 5
SNR $<$ 4 Minimal Mask 5 5 4 SNR $<$ 13 Minimal Mask 4 4
 [^2] The Parallel Red band uses the same mask sizes as the Scan Map Red band.
Figure 5: Examples of the six different mask sizes used in the Scan Map Blue dataset. For each example, the left column shows the cutout image with the target object centered, along with the corresponding mask radius and the SNR value of the central object. The right column displays the binary mask used for inpainting. Note that multiple objects may appear within a cutout, but the SNR value and mask radius are determined based on the source located at the center of the image.
In these masks, white pixels indicate areas considered as valid background, while the black region represents the area to be modified. These masks, while not being overly strict, guide the model in differentiating between regions that are likely to be part of the compact source and those that constitute valid background.
2.4.5 Input and Target Pairs
The approach is straightforward: we start with the entire field containing the injected sources, the corresponding target maps — these are the original images used to inject the sources — and a binary mask with the same size as the other two maps, incorporating the circular masks. For each coordinate, we extract smaller cutouts from all three maps: the image with the injected source centered, the corresponding binary mask, and the original image. The input to the network consists of the cutout of the image with the injected source at its center, along with the corresponding binary mask, while the cutout from the original image serves as the target. These input-mask-target pairs are used to train the network to learn how to remove the PSF and accurately estimate the background behind the source.
3 Methods
We developed a neural network architecture grounded in an inpainting approach to address the challenge of compact source removal and extraction. Inpainting refers to a technique used to reconstruct missing or masked parts of an image based on the surrounding information. For our method, we use dynamically constructed masks with the SNR values, as described in Section 2.4.4.
3.1 Architecture
The U-Net architecture by Ronneberger et al. (2015), originally designed for biomedical image segmentation, forms the backbone of our model. This architecture is characterized by its symmetric structure, comprising an encoder for downsampling and decoder for upsampling, connected by skip connections. This design is particularly effective in preserving spatial information throughout the network, a crucial aspect for our task to access background regions directly.
Building upon the foundational concept of U-Net, we also drew inspiration from the work of Isola et al. (2017), where they demonstrated the potential of such architecture in image-to-image translation tasks through their Conditional Generative Adversarial Network (cGAN). Their methodology highlighted the effectiveness of U-Net in learning complex, conditional mappings between input and output image domains.
We incorporated the Convolution-BatchNorm-ReLU module proposed by Radford et al. (2016) in their DCGAN network. This configuration, which includes a sequence of convolutional layers followed by batch normalization (Ioffe & Szegedy (2015)) and Rectified Linear Unit (ReLU) activation (Nair & Hinton (2010)), is known for its efficiency in deep convolutional networks. This is what Isola et al. (2017) used in their network as well, however, in contrast to the configuration in their generator network, we replaced the traditional convolutional layers with partial convolutional layers (PConv) [^3] https://github.com/naoto0804/pytorch-inpainting-with-partial-conv:
- We used and adapted this implementation for partial convolution.
, used in Liu et al. (2018). These layers are great at handling irregular (non-rectangle) holes in inpainting tasks, allowing the network to focus only on valid pixels in the convolutional process, and the mask is also updated layer-by-layer.
We opted for strided convolutions instead of traditional pooling layers for both downsampling and upsampling. This choice is guided by the aim to enable the network to learn its own spatial transformations (Springenberg et al., 2015). Differing from Liu et al. (2018), where nearest neighbor upsampling was used, we implemented transposed partial convolutional layers for detailed spatial reconstruction. These layers enable the network to autonomously learn upsampling, providing flexibility and accuracy in handling the complex spatial structures of astronomical data.
3.2 Implementation Details
3.2.1 Network Architecture
Our network, as depicted in Fig. 7, consists of an encoder and a decoder segment. The network takes the input image along with the corresponding mask. This combined input then passes through the network, ultimately resulting in a one-channel grayscale image, where any compact source(s) present in the input are removed. A key element in both segments is the Convolution-BatchNorm-ReLU (CBR) block, which sequentially applies partial convolution, batch normalization, and an activation function. To construct our model, we defined two main types of these blocks, specifically designed for the encoder and the decoder.
Figure 6: Core Components of the Architecture: The PCDown Block (left) downsamples and extracts features in the encoder, while the PCUp Block (right) upsamples in the decoder.
Figure 7: Illustration of our architecture: This U-Net-like architecture comprises 11 PCDown Blocks in the encoder segment and an additional 11 PCUp Blocks in the decoder segment. The corresponding feature vectors at layer $i$ in the encoder are copied and concatenated to the input of layer $n-i$ in the decoder, where $n$ is the total number of layers. This structure helps the network in accessing specific regions of the input image directly. Further details of the architecture and its parameters can be found in Table 4.
In the encoder part, our Partial Convolution Downsample (PCDown) block configuration (left side in Fig. 6) is constructed as follows:
  • Partial Convolution layer (Liu et al., 2018) is used to handle the inpainting task, specifically to focus only on valid (background) regions of the input image during the convolution process. It dynamically updates the mask after each layer, allowing efficient feature extraction.
  • Batch Normalization (Ioffe & Szegedy, 2015) is applied after convolution, which stabilizes the learning process by normalizing the output of the previous layer.
  • ReLU Activation Function (Nair & Hinton, 2010) is employed to introduce non-linearity into the network, enabling it to learn more complex patterns in the data.
The Partial Convolution Upsample (PCUp) block (right side in Fig. 6) in the decoder segment is similar but with some differences:
  • Transposed Partial Convolution serves the purpose of upsampling the feature maps. It combines the idea of transposed convolution (Long et al., 2015) with partial convolution.
  • Batch Normalization and Activation Functions are similar to those in the encoder, however, we used LeakyReLU (Maas et al., 2013; Xu et al., 2015) instead of ReLU with a negative slope parameter of 0.2.
3.2.2 Encoder
The encoder segment utilizes an alternating sequence of two block types to progressively downsample the input while extracting features:
  • Strided Convolution Blocks employ a PCDown block with a kernel size of 4, a stride of 2, and padding of 1. This configuration effectively reduces the spatial dimensions of the input image by half, enabling efficient downsampling by increasing the receptive field.
  • Feature Extraction Blocks follow each strided convolution block. These employ a PCDown block with a kernel size of 3, a stride of 1, and padding of 1, maintaining the spatial dimensions while further extracting detailed features from the downsampled inputs.
The channel dimensions in these layers start from the number of input channels (1 in our case) and increase progressively, reaching up to 512. The last layer of the encoder, PCDown11, differs from the usual strided convolution blocks in order to reduce the spatial dimensions to $1\times 1$ for $96\times 96$ pixel input images, forming a compact encoded feature vector that captures the essential characteristics of the input. See the details of the encoder architecture in Table 4.
Table 4: Details of Encoder and Decoder Architecture for 96x96 pixel inputs.
Encoder Layers Decoder Layers
Layer In Ch. Out Ch. Kernel Stride Padding Layer In Ch. Out Ch. Kernel Stride Padding
PCDown1 1 32 3 1 1 PCUp1 512 512 3 1 0
PCDown2 32 64 4 2 1 PCUp2 1024 512 4 2 1
PCDown3 64 64 3 1 1 PCUp3 1024 512 3 1 1
PCDown4 64 128 4 2 1 PCUp4 1024 256 4 2 1
PCDown5 128 128 3 1 1 PCUp5 512 256 3 1 1
PCDown6 128 256 4 2 1 PCUp6 512 128 4 2 1
PCDown7 256 256 3 1 1 PCUp7 256 128 3 1 1
PCDown8 256 512 4 2 1 PCUp8 256 64 4 2 1
PCDown9 512 512 3 1 1 PCUp9 128 64 3 1 1
PCDown10 512 512 4 2 1 PCUp10 128 32 4 2 1
PCDown11 512 512 3 1 0 PCUp11 64 1 3 1 1
 [^4] The encoder consists of 11 PCDown blocks, each employing a ReLU activation function. The decoder is composed of 11 PCUp blocks, using leaky ReLU activation function for all but the final block.
3.2.3 Decoder
The decoder segment of our network is designed to mirror the encoder’s structure in terms of the spatial dimensionality of the feature maps (see Fig. 7). This symmetry is crucial for the integration of skip connections, which connect each layer $i$ in the encoder and layer $n-i$ in the decoder, where $n$ is the total number of layers. Skip connections play an important role in preserving input details (Ronneberger et al., 2015). In the decoder (see Table 4), each layer is designed to match the parameters of the corresponding encoding layer. It comprises an alternating sequence of ”upsampling blocks” and ”feature refinement blocks”:
  • Upsampling Blocks utilize PCUp blocks, where the transposed partial convolution is configured with a kernel size of 4, a stride of 2, and a padding of 1. This setup incrementally increases the spatial dimensions of the feature maps, effectively doubling them at each step.
  • Feature Refinement Blocks follow each upsampling block, a PCUp block with a kernel size of 3, a stride of 1, and padding of 1 is employed. These blocks do not change the spatial dimensions but instead focus on refining and enriching the features that have been upsampled.
Additionally, as shown in the table, the number of input channels in the decoder is doubled due to the skip connections, where the output of the corresponding encoder layer is combined with the input of the matching decoder layer. This allows the decoder to have direct access to spatial information from earlier stages of the network, improving feature recovery and spatial consistency. The first layer of the decoder, PCUp1, diverges from the standard ”Upsampling” and ”Feature Refinement” block design to mirror the last block of the encoder, where we employed different parameters to compress the $96\times 96$ pixel input into a $1\times 1$ spatial dimension feature vector.
3.3 Loss
The loss function used during training plays a crucial role in guiding the model to produce accurate source-free images. Our loss function is an adaptation of the one proposed in Liu et al. (2018), with some modifications customized to our specific requirements. We added the loss terms one-by-one and closely examined how each affected learning on a smaller subset of our dataset. In the following subsections, we describe the final loss terms used after our experiments.
3.3.1 L1 Loss for Hole and Valid Regions
Following the approach in Liu et al. (2018), we compute the L1 loss separately for the hole (masked) and valid (unmasked) regions of the image. The L1 losses for masked and unmasked regions are defined mathematically as follows:
$\begin{split}\mathcal{L}_{\text{hole}}&=\frac{1}{N_{\mathbf{I}_{\mathrm{gt}}}}\left|\left|\left(1-\mathbf{M}\right)\odot\left(\mathbf{I}_{\text{out}}-\mathbf{I}_{\text{gt}}\right)\right|\right|_{1}\\ \mathcal{L}_{\text{valid}}&=\frac{1}{N_{\mathbf{I}_{\mathrm{gt}}}}\left|\left|\mathbf{M}\odot\left(\mathbf{I}_{\text{out}}-\mathbf{I}_{\text{gt}}\right)\right|\right|_{1},\end{split}$ (2)
where $\mathbf{I}_{\mathrm{out}}$ is the output image generated by the network (the estimated background image), $\mathbf{I}_{\mathrm{gt}}$ is the ground truth image (the original background), and $\mathbf{M}$ represents the mask, with element-wise multiplication denoted by $\odot$. $N_{\mathbf{I}_{\mathrm{gt}}}$ denotes the total number of pixel values in an image, calculated as $C\times H\times W$, where $C$ is the channel size, and $H$ and $W$ are the height and width of the image $\mathbf{I}_{\text{gt}}$, respectively. $\mathcal{L}_{\text{hole}}$ targets the hole regions, where the sources are located, while $\mathcal{L}_{\text{valid}}$ ensures the integrity of the existing background is preserved.
3.3.2 Style Loss
The style loss, introduced by Gatys et al. (2015), and utilized in the context of inpainting by Liu et al. (2018), is also an integral part of our loss function. It helps in capturing the texture and style of the image, ensuring that the synthesized regions (background behind the source) blend seamlessly with the surrounding (valid) areas. The style loss is formulated as follows:
$\begin{split}\mathcal{L}_{\text{style}_{\text{out}}}&=\sum_{p=0}^{P-1}\frac{1}{C_{p}C_{p}}\left|\left|K_{p}\left(\left(\Psi_{p}^{\mathbf{I}_{\text{out}}}\right)^{\top}\left(\Psi_{p}^{\mathbf{I}_{\text{out}}}\right)-\left(\Psi_{p}^{\mathbf{I}_{\text{gt}}}\right)^{\top}\left(\Psi_{p}^{\mathbf{I}_{\text{gt}}}\right)\right)\right|\right|_{1}\\ \mathcal{L}_{\text{style}_{\text{comp}}}&=\sum_{p=0}^{P-1}\frac{1}{C_{p}C_{p}}\left|\left|K_{p}\left(\left(\Psi_{p}^{\mathbf{I}_{\text{comp}}}\right)^{\top}\left(\Psi_{p}^{\mathbf{I}_{\text{comp}}}\right)-\left(\Psi_{p}^{\mathbf{I}_{\text{gt}}}\right)^{\top}\left(\Psi_{p}^{\mathbf{I}_{\text{gt}}}\right)\right)\right|\right|_{1}\end{split}$ (3)
The Gram matrix, $\left(\Psi_{p}^{\mathbf{I}}\right)^{\top}\left(\Psi_{p}^{\mathbf{I}}\right)$, captures the correlation between different filter responses, effectively representing the style of the astronomical image by comparing their feature representations as extracted by a pre-trained convolutional neural network, like VGG-16 (Simonyan & Zisserman, 2015), where $\Psi_{p}$ is the feature map at layer $p$. The loss is calculated separately for the output image and a composite image ($\mathbf{I}_{\text{comp}}$), which is the raw output image, but with the non-hole pixels directly set to ground truth. The style loss then measures the difference in these correlations between the generated, source-free image ($\mathbf{I}_{\text{out}}$ or $\mathbf{I}_{\text{comp}}$) and the ground truth ($\mathbf{I}_{\text{gt}}$). $C_{p}$ is the number of channels in the feature map at layer $p$, and $K_{p}$ is a normalization factor.
3.3.3 Structural Similarity Index Measure (SSIM) Loss
We incorporate the Structural Similarity Index Measure (SSIM) loss (Wang et al. (2004)) in our loss function, replacing the perceptual and total variation losses used in Liu et al. (2018), as they did not improve background estimation and blending in our setting. The SSIM loss is advantageous in preserving the structural information in images, which is important when maintaining the integrity of dust cloud structures in such images. The SSIM loss is defined as:
$\mathcal{L}_{\text{SSIM}}=1-\text{SSIM}(\mathbf{I}_{\text{gt}},\mathbf{I}_{\text{pred}}),$ (4)
where SSIM is the structural similarity index measure between the ground truth and the predicted image, and is defined as:
$\text{SSIM}(x,y)=\frac{\left(2\mu_{x}\mu_{y}+c_{1}\right)\left(2\sigma_{xy}+c_{2}\right)}{\left(\mu_{x}^{2}+\mu_{y}^{2}+c_{1}\right)\left(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2}\right)},$ (5)
where $\mu_{x}$ and $\mu_{y}$ are the pixel sample means, $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the variances of the images, $\sigma_{xy}$ is the covariance of the two image, $c_{1}=\left(k_{1}L\right)^{2}$ and $c_{2}=\left(k_{2}L\right)^{2}$ are two variables to stabilize the division with weak denominator and $L$ is the dynamic range of the pixel-values; $k_{1}=0.01$ and $k_{2}=0.03$ by default. We used the implementation from the Kornia library (E. Riba et al., 2020), a differentiable computer vision library for PyTorch that provides a wide range of tools seamlessly integrated into deep learning models.
3.3.4 Source Loss
In addition to the previously described components, we included another loss to directly compare the original source and the extracted source, which we call the ”source loss”. This comparison ensures accurate extraction of the flux density and reconstruction of the PSF. The source loss can be expressed mathematically as follows:
$\mathcal{L}_{\text{source}}=\left|\left|\underbrace{\left(\mathbf{I}_{\text{orig}}-\mathbf{I}_{\text{gt}}\right)}_{\text{real source}}-\underbrace{\left(\mathbf{I}_{\text{orig}}-\mathbf{I}_{\text{pred}}\right)}_{\text{extracted source}}\right|\right|_{1},$ (6)
where $\mathbf{I}_{\text{orig}}$ represents the original input image, $\mathbf{I}_{\text{gt}}$ is the ground truth sourceless image, and $\mathbf{I}_{\text{pred}}$ is the predicted source-free image. The difference between $\mathbf{I}_{\text{orig}}$ and $\mathbf{I}_{\text{gt}}$ reveals the real injected PSF (with the scaled flux values and rotated orientation), while the difference between $\mathbf{I}_{\text{orig}}$ and $\mathbf{I}_{\text{pred}}$ represents the source as extracted by the network, see the example in Fig. 8.
Figure 8: Visualization of the source loss. The top row displays the input image, the ground truth target (source-less image), and the output prediction (during the training process), respectively. The bottom row illustrates the real source (difference between the original image and the ground truth sourceless image), extracted source (difference between the original image and the predicted source-free image), and the difference between these two sources.
3.3.5 Final Loss Function
The final loss function for our network is a weighted sum of the individual loss components. We initially used the weights determined by Liu et al. (2018) for the components we also utilize, as they worked well. For the new components, we set their weights so that their values were of a similar magnitude to the others, ensuring they had a comparable impact. These values were calculated over a few epochs using different batches. For our final model, we used the following weighting:
$\mathcal{L}_{\text{total}}=2\mathcal{L}_{\text{valid}}+6\mathcal{L}_{\text{hole}}+120(\mathcal{L}_{\text{style}_{\text{out}}}+\mathcal{L}_{\text{style}_{\text{comp}}})+0.4\mathcal{L}_{\text{SSIM}}+\mathcal{L}_{\text{source}}.$ (7)
3.4 Astropy and Aperture Photometry
Astropy is a widely used open-source Python library for astronomy that provides a comprehensive set of tools for data manipulation, analysis, and visualization (Astropy Collaboration et al., 2013, 2018, 2022).
In this study, we used Astropy’s aperture_photometry function (Bradley et al., 2024), which is designed to measure flux densities in defined regions of images. The aperture_photometry function allows us to specify a circular aperture around a source, within which it sums the pixel values to estimate the source’s flux.
We also performed aperture photometry correction to account for the portion of the PSF that lies outside the aperture. This correction is important for obtaining a more accurate flux measurement, as parts of the PSF of the sources fall outside the defined aperture radius.
We applied corrected aperture photometry in two distinct ways. First, on the original image, we used a circular annulus around each source to estimate the mean background level, which we subtracted from the aperture photometry result to obtain the source flux. Then, to assess our model’s performance, we measured the flux on the image where the source was extracted by our network. For this comparison, we performed aperture photometry with correction but without subtracting a background value, as our model estimates the background directly by removing the source.
3.5 Training Procedure
3.5.1 Setup and Hyperparameters
The training inputs were 1-channel images of size $96\times 96$ pixels, as detailed in Section 2.4.5. The model was trained using a batch size of 64.
For the optimization process, we employed the Adam optimizer (Kingma & Ba (2017)), known for its efficiency in handling sparse gradients and adaptive learning rates. The initial learning rate was set at 0.005. To enhance the training process and adapt the learning rate based on validation loss performance, we utilized the ReduceLROnPlateau scheduler (Paszke et al. (2019)) with a patience of 5 epochs and a reduction factor of 0.5. This approach helps in fine-tuning the learning rate, ensuring the model converges to an optimal solution efficiently.
3.5.2 Epochs and Validation
The training of our model followed a systematic approach, after each epoch of training on the training dataset, a validation loop was conducted on the separate validation dataset. During these validation loops, we closely observed the performance of the model using the final loss function (Equation 7).
The learning rate adjustments were managed by the ReduceLROnPlateau scheduler, which was triggered based on the final loss observed on the validation dataset. This ensured that learning rate modifications were aligned with the overall performance improvements of the model.
To further refine the model, we implemented a checkpointing system where the model’s parameters were saved each time a new lowest value of the final loss was achieved on the validation dataset, mitigating issues related to overfitting or training stagnation.
4 Results
Our main result is the open-source tool we developed and made available for the community to use in removing compact sources, hosted at our GitHub repository  [^5] https://github.com/xMediMadix/compact-source-removal. Users can employ the pre-trained model for source removal from Herschel images, or re-train the model on custom datasets. The repository includes three main configuration files, each designed for one of the tasks described below, along with a detailed README.md file explaining how to utilize the code effectively. In Section 5, we will evaluate our model’s performance across various aspects. This includes an accuracy comparison of the photometry pipeline we used on standard star models, a comparison between the measured flux densities of extracted sources and the known values of the injected PSFs, and an analysis of how the model impacts the power spectrum and the pixel intensity distributions of the maps.
4.1 Removal on Simulation
We provide resources to demonstrate our model’s evaluation on simulations, including examples to guide users through the process. The tool iterates through simulation files, applies source removal using the trained model, and exports the resulting source-free background images, along with .csv files that contain photometric results for each source at specified aperture sizes.
Figure 9 illustrates an example of this process. In this simulation, 90 sources with flux densities randomly chosen between 40 and 10,000 mJy were injected at random positions within the Scan Map Blue observation data with oid=1342237154. The top panel shows the simulated image with the injected sources, while the bottom panel displays the output of our evaluation tool, with compact sources removed to reveal the extended background emission.
Figure 9: Example of simulated source removal. Top: Simulated image with 90 injected sources (randomly chosen flux densities from 40 to 10,000 mJy) within the Scan Map Blue observation data (oid=1342237154). Bottom: Model output with sources removed, providing a source-free background image.
4.2 Removal on Real Data
This core functionality is likely the most beneficial for the community. Users can specify parameters directly in a configuration file, defining aspects such as Observing Mode, band, aperture sizes for photometry, and other settings. We include examples to assist users. The tool processes input images by iterating through the coordinates of compact sources, removing them, and exporting a map with the sources removed, along with a .csv file that reports photometric results for the specified apertures. Users can input their own coordinates or use the new Herschel/PACS Point Source Catalog created by Marton et al. (2024).
Figure 10 shows examples of this functionality applied to real data. The top row depicts an original and processed image from the Scan Map Green observation with oid=1342185553. The left image in the first row shows the original input, and the right image displays the result after our tool has removed compact sources, leaving a background-only map. Similarly, the bottom row shows another example from Scan Map Green observation oid=1342213500. In both examples, the coordinates of compact sources were taken from the catalog by Marton et al. (2024). These examples demonstrate the tool’s effectiveness in removing cataloged compact sources from real Herschel data, allowing users to analyze the underlying extended emission without interference from compact sources.
Figure 10: Examples of compact source removal on real data. Top row: Original (left) and processed (right) Scan Map Green image (oid=1342185553). Bottom row: Original (left) and processed (right) Scan Map Green image (oid=1342213500). Coordinates were taken from the new Herschel/PACS Point Source Catalog.
4.3 Re-training on Custom Data
While we include trained models for both Scan Map and Parallel Observing Modes, we also provide a small subset of our training data, enabling users to run the training pipeline and gain insight into its operation. The tool has been designed to allow for the inclusion of custom data sets, allowing users to implement their own data and train new models using our pipeline if required.
5 Discussion
In this section, we discuss our findings by presenting evaluation results of the model’s performance across different aspects. To ensure the reliability of our approach, we tested the model not only on simulated datasets but also on real sources by comparing the extracted fluxes with those of standard star models. This helped us confirm the accuracy of the extracted fluxes against well-calibrated reference values. We also analyzed the photometric accuracy on simulations to evaluate how well the model reconstructed the background and extracted sources. Additionally, we further assessed the preservation of image properties by evaluating the power spectrum and pixel intensity distribution of the resulting images. These evaluations provide a thorough overview of the model’s strengths and areas for improvement.
5.1 Comparison with Standard Star Models
Studies by Balog et al. (2014) and Klaas et al. (2018) have previously examined the flux calibration of the PACS photometer. We applied our network to estimate the background and extract the sources, then performed aperture photometry as discussed in Section 3.4 to measure the flux density of selected standard stars. We used the fluxes provided in these studies as references to evaluate both the accuracy of the aperture_photometry function and the source removal performance of our network. Table 5 summarizes the comparison of the reference flux densities from the studies with the measured flux densities of the sources extracted by our network across the $70\mu\mathrm{m}$, $100\mu\mathrm{m}$, and $160\mu\mathrm{m}$ bands.
These results validate that we can reliably use aperture_photometry on the sources extracted by our network. Additionally, the extracted fluxes closely match the reference flux densities of the standard stars, as demonstrated by the flux ratios we recover in Table 5. This confirms the ability of our model to accurately reconstruct real compact sources, ensuring minimal deviation from the established calibration values.
Some values are missing from the table due to limitations in the data retrieval process. We used the coordinates of the standard stars to download the maps from the Herschel Science Archive website, performed a centroid fitting to refine the source positions, and then evaluated our model at those locations. However, in some cases, the calibration measurements were not processed up to at least Level2.5, preventing us from including them in our analysis. For the $160\mu\mathrm{m}$ red band, some values at the bottom of the table are also missing because the sources were below the detection limit at that wavelength, making accurate flux measurements impossible.
Table 5: Comparison of PACS standard star reference flux densities with measured flux densities from extracted sources.
Scan Map Blue - 70 $\mu\mathrm{m}$ Scan Map Green - 100 $\mu\mathrm{m}$ Scan Map Red - 160 $\mu\mathrm{m}$
name ra dec ref. ref. err. ours ratio ref. ref. err. ours ratio ref. ref. err. ours ratio src.
[deg] [deg] [mJy] [mJy] [mJy] [mJy] [mJy] [mJy] [mJy] [mJy] [mJy]
$\alpha$ Boo 213.915 19.182 15,434.0 15,383.9 0.997 7,509.0 7,506.1 1.000 2,891.0 2,900.5 1.003 (1)
$\alpha$ Tau 68.980 16.509 14,131.0 14,098.7 0.998 6,909.0 6,882.7 0.996 2,677.0 2,677.5 1.000 (1)
$\beta$ And 17.433 35.621 5,594.0 5,608.5 1.003 2,737.0 2,751.2 1.005 1,062.0 1,068.0 1.006 (1)
$\alpha$ Cet 45.570 4.090 4,889.0 4,884.9 0.999 2,393.0 2,365.2 0.988 928.0 925.0 0.997 (1)
$\gamma$ Dra 269.152 51.489 3,283.0 3,283.4 1.000 1,604.0 1,607.8 1.002 621.0 623.5 1.004 (1)
$\beta$ Gem 116.329 28.026 2,457.0 140.8 2,634.1 1.072 1,190.0 68.2 1,226.7 1.031 455.9 26.1 468.7 1.028 (2)
$\alpha$ Ari 31.793 23.462 1,707.0 100.7 1,705.6 0.999 831.4 49.1 833.5 1.003 321.0 18.9 323.1 1.006 (2)
$\epsilon$ Lep 76.365 -22.371 1,182.0 69.7 1,185.0 1.003 576.2 34.0 578.6 1.004 222.7 13.1 213.8 0.960 (2)
$\omega$ Cap 312.955 -26.919 857.7 51.7 851.6 0.993 418.0 25.2 412.8 0.988 161.5 9.7 159.9 0.990 (2)
$\eta$ Dra 245.998 61.514 479.5 16.2 515.8 1.076 232.6 8.0 220.7 0.949 89.4 3.1 83.0 0.928 (2)
$\delta$ Dra 288.139 67.662 428.9 24.5 436.9 1.019 207.7 11.8 210.9 1.015 79.6 4.5 85.0 1.067 (2)
$\theta$ Umi 232.854 77.349 286.2 16.2 288.5 1.008 139.5 7.9 53.9 3.1 (2)
HD 41047 90.318 -33.912 195.6 11.7 95.4 5.7 95.7 1.003 36.9 2.2 28.2 0.764 (2)
42 Dra 276.496 65.563 153.7 4.6 149.1 0.970 75.3 2.3 72.1 0.957 29.4 0.9 30.1 1.023 (2)
HD 138265 231.964 60.670 115.9 4.0 113.8 0.982 56.8 2.0 55.8 0.982 22.2 0.8 20.5 0.925 (2)
HD 159330 262.682 57.877 64.2 2.1 62.2 0.969 31.5 1.0 31.3 0.995 12.3 0.4 12.7 1.033 (2)
HD 152222 251.769 67.267 39.4 1.9 40.7 1.034 19.3 1.0 20.6 1.065 7.5 0.1 (2)
HD 39608 87.402 -60.676 30.9 1.2 30.6 0.990 15.1 0.6 5.9 0.2 (2)
$\delta$ Hyi 35.437 -68.659 22.9 0.8 20.9 0.914 11.2 0.4 10.3 0.920 4.4 0.2 (2)
 [^6] Columns include star name, J2000 coordinates, reference and measured flux densities at $70\mu\mathrm{m}$, $100\mu\mathrm{m}$, and $160\mu\mathrm{m}$ with errors (when available), their ratios, and the source of reference values. References are (1) Balog et al. (2014) and (2) Klaas et al. (2018).
5.2 Photometric Accuracy
As described in Section 2.3.2, we preserved several fields for testing purposes, with the numbers shown in Table 2. These fields were not used in either the training or validation loops, making them suitable for testing our model on unseen cases.
An important aspect of evaluating our model’s performance involves measuring the flux densities of the extracted sources and comparing them to the theoretical flux densities of the injected PSFs. If the measured flux density of the extracted source closely aligns with the known value, it indicates that our model has accurately estimated the background and successfully removed the source with precision.
We applied the same processing steps to our test dataset as we did to the other datasets: each image was per-image normalized, mask sizes were determined using the same parameters, and all inputs were passed through our pre-trained model to create source-free, background-only images for all injected sources. We obtained the extracted source by subtracting the model’s predicted background from the original input image. Next, we performed aperture photometry on this background-subtracted source with various aperture sizes, ranging from 4 to 18 arcseconds, to measure the extracted flux density. We also performed circular annulus background estimation on the input images, as described in Section 3.4, and compared these flux density values to the known injected flux densities by calculating their ratios.
Figure 11 shows the results for the Scan Map mode, while Fig. 12 illustrates results for the Parallel mode. In these figures, blue box plots represent the ratios of the flux densities measured on sources extracted by our model to the known injected flux densities. Additionally, gray box plots show the flux ratios measured directly from the input images using circular annulus background estimation, without model-based source removal. The horizontal line at $y=1$ indicates the ideal flux ratio, where the measured flux exactly matches the known value. While the gray boxes reflect the flux density estimations from standard photometry, the blue boxes assess the model’s effectiveness in accurately estimating the background by removing the injected source. A flux ratio of 1 indicates that the model has precisely removed the injected flux, achieving an accurate background estimate.
Two key results are visible in the figures. First, our model accurately removes the sources with high precision, as seen in the close alignment of the green median lines in the blue boxes with the ideal line. Second, not only is the source removal accurate on average, but for most cases, the flux estimates for the extracted sources are more consistent than those obtained with circular annulus background estimation, as shown in the narrower blue box plots. This reduced interquartile range (IQR) suggests that the model offers a more stable and reliable flux measurements with less variation around the median compared to the annulus method.
It is worth noting that the IQR is generally larger for sources with lower flux densities. This is because we set the lowest flux levels for all fields where sources only begin to appear against the background, leading to larger deviations from the ideal as these flux levels approach the detection threshold. Also, in the Scan Map $70\mu\mathrm{m}$ (BS) band plot, starting at around $13,000$ mJy, a noticeable increase in deviation can be observed, which is due to sources in regions with particularly complex backgrounds, such as dense clouds and complex structures. Despite these challenges, the model-based approach generally provides more accurate background estimates than the circular annulus method in these difficult environments.
Figure 11: Photometric accuracy evaluation for the Scan Map mode across three bands. Top: Scan Map blue ($70\mu\mathrm{m}$ BS band); Middle: Scan Map green ($100\mu\mathrm{m}$ BL band); Bottom: Scan Map red ($160\mu\mathrm{m}$ R band). Blue boxes represent the flux density ratios of sources extracted by our model to the known injected fluxes, while gray boxes show ratios using standard circular annulus background estimation without source removal. The line at $y=1$ indicates perfect flux matching.
Figure 12: Photometric accuracy evaluation for the Parallel mode across three bands. Top: Parallel blue ($70\mu\mathrm{m}$ BS band); Middle: Parallel green ($100\mu\mathrm{m}$ BL band); Bottom: Parallel red ($160\mu\mathrm{m}$ R band).
5.3 Image comparison
To evaluate the performance of our method, we further compare the properties of the original source-free images with the source-free predictions obtained after removing simulated sources. In this subsection, we present two analyses to assess the alignment between these maps: the power spectrum and the pixel intensity distribution. The aim of these approaches is to ensure that the predicted maps preserve the structural and statistical characteristics of the original source-free images, demonstrating the effectiveness of the source removal process.
The power spectrum analysis examines the spatial frequency domain, providing insights into whether the model has preserved the underlying structures and avoided introducing artifacts. Meanwhile, the pixel intensity distribution analysis evaluates the statistical properties of pixel values in the predicted maps, ensuring that the overall image characteristics remain consistent with the original. Together, these evaluations provide a great assessment of the method’s ability to accurately reproduce the original source-free image properties.
Figure 13: Left column: Direction independent Fourier transform power spectra of 3 different fields before adding artificial sources (blue line), with the artificial sources (orange) and after their removal (green). Power spectra are normalized. Middle column: Difference of the power spectrum before adding sources and after their removal. Right column: Probability density functions before adding artificial sources and after their removal. Blue histograms show the PDF of the unmodified map, red histograms show the PDFs of the predicted maps.
5.3.1 Power spectrum
The Fourier transform power spectra of the different maps provide useful insight into the properties of the signal introduced by our tool. In Fig. 13 we demonstrate a comparison with one case in each filter band. These fields were chosen from the test dataset, and for each field, we generated random positions for the sources. For each random position, a random flux density and orientation were also assigned, ensuring that the fields represent the most diverse and variable cases possible. These maps are similar to the example shown in Fig. 9.
The images used for this comparison are as follows: Scan Map Blue oid=1342216409, Scan Map Green oid=1342244224, and Scan Map Red oid=1342249136. This order corresponds to the arrangement of the bands from top to bottom in Fig. 13.
Plots in the left hand column show the power spectra of the fields. They clearly illustrate that adding the artificial sources significantly modifies the power spectrum (orange lines), especially at the spatial scale similar to the size of the point sources. They also clearly show that the original (blue lines) and the predicted maps (green lines) are very close and actually cannot be distinguished on these plots, the plotting lines do completely overlap. To highlight any subtle differences, we also plotted the difference between the power spectra of the original and predicted maps, shown in the middle column. These plots reveal that the differences in the normalized spectra are only on the order of 10-4 to 10-10, depending on the spatial frequency.
5.3.2 Pixel intensity distribution
As an additional test, we examined the PDF of the fields. Examples are shown in the right column of Fig. 13, focusing on the masked regions where compact sources were injected and subsequently removed. The PDFs of the original source-free maps (blue histograms) and the predicted source-free maps (red histograms) show strong agreement. This demonstrates that the model accurately reconstructs the extended emission and preserves the statistical properties of the pixel intensity distribution, including noise features and background variations.
Table LABEL:table:metrics_complete provides a detailed comparison of pixel intensity statistics in the masked regions. Metrics such as the mean, median, standard deviation, skewness, and interquartile range were analyzed. The alignment of mean and median values between the original and predicted maps indicates that the model effectively recovers the background emission without introducing biases in brightness. Furthermore, the agreement in standard deviation and IQR confirms that the model preserves the spread and variability of the background intensity. The consistency in skewness further highlights the model’s ability to maintain asymmetry in the pixel intensity distribution, reflecting complex structures in the interstellar medium.
These results validate that the model not only removes compact sources across a range of flux densities but also preserves the underlying extended emission and statistical integrity of the images. This consistency across all test fields, observing modes, and filter bands delivers reliable maps for subsequent astrophysical analysis.
6 Summary
We introduced a deep learning approach for effective source removal from photometric observations. The goal of the method was two-fold: analysis of the extended emission is less effective in the presence of point sources, as their presence is disturbing to the power spectrum and other image analysis techniques. On the other hand, the fluctuating background is a problem for many photometry methods, as the background cannot be easily estimated.
To address these challenges, we presented a novel source removal method tailored for photometric observations made by the Herschel/PACS instrument. This approach was based on a U-Net-like architecture enhanced by Partial Convolution blocks. It integrated carefully selected loss functions, including L1 loss, style loss, SSIM loss, and custom source loss, ensuring seamless background reconstruction and accurate source extraction. The use of encoder-decoder symmetry with skip connections effectively preserved spatial information while extracting complex features from the input data. A dynamically calculated mask further improved the efficiency and accuracy of the process. Key design elements included the Partial Convolution Downsample and Upsample blocks, tailored for handling masked inputs and achieving efficient inpainting of source-free regions.
The performance of the architecture was evaluated using multiple approaches. First, we assessed background reconstruction in the masked areas by comparing the amount of injected and removed flux, demonstrating that our method provided more consistent results for aperture photometry compared to the standard approach, which lacked robust background estimation. We also compared the properties of the original and predicted images through Fourier transform power spectra and probability density functions, finding that the original image characteristics were preserved across all spatial scales and intensity distributions. Additionally, we validated the accuracy of our method on real sources by comparing extracted fluxes to reference values from standard star models, confirming its reliability on real data.
In conclusion, the proposed approach offers a powerful tool for the community, significantly reducing manual efforts in image processing while enhancing the reliability of background estimations and source extractions. Although the method was applied to Herschel/PACS observations, it can be trained on other photometric maps as well. We also provided a Python package with tutorials and examples for the users.
Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation program under grant number 101004141. G.M. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 521 selected in 2021, Revisiting Star Formation in the Era of Big Data (https://teams.issibern.ch/starformation/).
Appendix A Details of the Simulations
Table 6: Details of the dataset for the Scan Map Observing Mode in the Blue band.
Field # Sources Min Flux Max Flux Dataset Total Sources # Simulations
1342227809_L2_5 190 700 30000 Training 5890 31
1342195485_L2_5 170 20 6000 Training 5100 30
1342216487_L2_5 170 150 15000 Training 5610 33
1342258816_L2_5 150 150 14000 Training 4800 32
1342191945_L2_5 145 10 5000 Training 4350 30
1342241633_L2_5 120 10 6000 Training 3720 31
1342188474_L2_5 100 10 3500 Training 2700 27
1342246718_L2_5 100 20 8000 Training 3200 32
1342191106_L2_5 90 2000 40000 Training 2520 28
1342183905_L3 70 10 4000 Training 1960 28
1342187003_L2_5 70 40 6000 Training 2030 29
1342191805_L2_5 70 80 14000 Training 2450 35
1342191801_L2_5 70 80 15000 Training 2520 36
1342187077_L2_5 60 60 6000 Training 1680 28
1342237428_L2_5 60 20 7000 Training 1860 31
1342217754_L2_5 55 40 8000 Training 1705 31
1342217752_L2_5 55 200 35000 Training 2200 40
1342217758_L2_5 55 250 30000 Training 2035 37
1342228952_L2_5 55 20 8000 Training 1760 32
1342190339_L2_5 50 10 4000 Training 1400 28
1342209668_L2_5 50 20 4000 Training 1350 27
1342227094_L2_5 50 10 6000 Training 1550 31
1342196757_L2_5 50 10 6000 Training 1550 31
1342226731_L2_5 50 40 8000 Training 1550 31
1342231674_L2_5 50 40 6000 Training 1450 29
1342205228_L3 50 900 40000 Training 1650 33
1342228913_L2_5 50 40 10000 Training 1650 33
1342227049_L2_5 50 100 11000 Training 1550 31
1342218780_L2_5 30 10 3000 Training 780 26
1342188204_L2_5 145 20 6000 Validation 4350 30
1342244881_L2_5 70 20 7000 Validation 2170 31
1342192767_L3 70 3500 40000 Validation 1750 25
1342191819_L2_5 55 400 35000 Validation 1980 36
1342217448_L2_5 55 150 14000 Validation 1760 32
1342216507_L2_5 50 10 3500 Validation 1350 27
1342245206_L3 45 10 6000 Validation 1395 31
1342249954_L2_5 200 100 12000 Test 6400 32
1342212042_L2_5 135 40 5500 Test 3780 28
1342216409_L2_5 115 40 3500 Test 2875 25
1342237154_L2_5 90 60 10000 Test 2880 32
1342213538_L3 40 20 3000 Test 1000 25
1342227047_L2_5 40 1750 40000 Test 1160 29
 [^7] The columns from left to right represent: the observation ID (field); the number of sources per simulation injected into each simulation for this field; minimum and maximum flux density levels; the resulting dataset split (Training, Validation, or Test); the total number of sources (calculated as the number of simulations multiplied by the number of sources per simulation); and finally, the number of simulations, which corresponds to the range of flux levels between the minimum and maximum flux densities.
Table 7: Details of the dataset for the Scan Map Observing Mode in the Green band.
Field # Sources Min Flux Max Flux Dataset Total Sources # Simulations
1342188880_L2_5 190 200 30000 Training 7220 38
1342216587_L2_5 185 40 7000 Training 5550 30
1342227811_L2_5 185 400 40000 Training 7030 38
1342212040_L2_5 170 20 12000 Training 6120 36
1342204085_L2_5 170 250 40000 Training 6970 41
1342188882_L2_5 140 150 30000 Training 5600 40
1342185555_L2_5 75 150 20000 Training 2700 36
1342241313_L3 65 20 20000 Training 2730 42
1342192771_L3 60 8000 40000 Training 1140 19
1342187005_L2_5 60 60 16000 Training 2280 38
1342237430_L2_5 55 40 18000 Training 2200 40
1342191821_L2_5 50 900 40000 Training 1650 33
1342191807_L2_5 50 200 22500 Training 1750 35
1342256983_L2_5 45 40 10000 Training 1485 33
1342228954_L2_5 45 80 22500 Training 1800 40
1342263882_L2_5 40 60 14000 Training 1440 36
1342203102_L2_5 40 60 15000 Training 1480 37
1342209035_L2_5 35 200 27500 Training 1295 37
1342207059_L2_5 35 150 25000 Training 1330 38
1342204298_L2_5 35 150 27500 Training 1365 39
1342193498_L2_5 30 40 8000 Training 930 31
1342269038_L2_5 30 80 14000 Training 1050 35
1342193548_L2_5 30 200 35000 Training 1200 40
1342193520_L2_5 30 150 35000 Training 1260 42
1342191815_L2_5 30 300 30000 Training 1080 36
1342270796_L2_5 30 100 18000 Training 1110 37
1342185545_L2_5 30 150 20000 Training 1080 36
1342259541_L2_5 25 20 6000 Training 750 30
1342231368_L2_5 25 60 8000 Training 750 30
1342218701_L2_5 185 60 8000 Validation 5550 30
1342241507_L2_5 100 10 13000 Validation 3800 38
1342267202_L3 55 80 25000 Validation 2255 41
1342261768_L2_5 45 40 9000 Validation 1440 32
1342231676_L2_5 45 60 20000 Validation 1800 40
1342193494_L2_5 30 150 27500 Validation 1170 39
1342230107_L2_5 30 150 20000 Validation 1080 36
1342237204_L2_5 135 40 8000 Test 4185 31
1342237206_L2_5 100 200 40000 Test 4200 42
1342188205_L2_5 100 150 10000 Test 2800 28
1342244224_L2_5 70 40 7000 Test 2100 30
1342189388_L2_5 40 80 14000 Test 1400 35
1342217764_L2_5 40 250 25000 Test 1400 35
 [^8] The columns are organized as detailed in Table 6.
Table 8: Details of the dataset for the Scan Map Observing Mode in the Red band.
Field # Sources Min Flux Max Flux Dataset Total Sources # Simulations
1342247320_L3 180 80 10000 Training 5580 31
1342228445_L2_5 150 150 15000 Training 4950 33
1342215726_L3 145 1750 40000 Training 4205 29
1342217454_L2_5 140 80 8000 Training 4060 29
1342221286_L2_5 140 200 20000 Training 4760 34
1342215595_L2_5 110 150 16000 Training 3740 34
1342220089_L2_5 90 150 18000 Training 3150 35
1342188204_L3 85 80 16000 Training 3145 37
1342258816_L2_5 85 500 40000 Training 3145 37
1342188882_L2_5 85 400 40000 Training 3230 38
1342204085_L2_5 85 700 40000 Training 2975 35
1342195485_L2_5 85 60 15000 Training 3145 37
1342212040_L3 85 60 9000 Training 2635 31
1342237206_L2_5 80 300 40000 Training 3200 40
1342224999_L2_5 65 200 15000 Training 2015 31
1342220308_L2_5 50 250 15000 Training 1500 30
1342220778_L2_5 50 150 12000 Training 1500 30
1342216401_L3 45 60 8000 Training 1350 30
1342216501_L2_5 40 80 10000 Training 1240 31
1342222157_L2_5 40 80 10000 Training 1240 31
1342216415_L2_5 40 80 10000 Training 1240 31
1342241507_L2_5 40 80 15000 Training 1440 36
1342220784_L2_5 35 100 7000 Training 945 27
1342249134_L2_5 165 600 40000 Validation 5940 36
1342220818_L2_5 90 150 15000 Validation 2970 33
1342219078_L2_5 90 150 25000 Validation 3420 38
1342218701_L2_5 60 100 15000 Validation 2100 35
1342249269_L3 130 300 40000 Test 5200 40
1342252015_L2_5 100 125 14000 Test 3300 33
1342223340_L2_5 90 150 16000 Test 3060 34
1342249136_L3 70 80 11000 Test 2240 32
1342204168_L2_5 50 80 7000 Test 1400 28
 [^9] The columns are organized as detailed in Table 6.
Table 9: Details of the dataset for the Parallel Observing Mode across all bands.
Band Field # Sources Min Flux Max Flux Dataset Total Sources # Simulations
Blue 1342213208_L2_5 175 200 30000 Training 6650 38
Blue 1342205093_L2_5 155 600 40000 Training 5580 36
Blue 1342214504_L2_5 130 200 40000 Training 5460 42
Blue 1342206679_L2_5 120 200 30000 Training 4560 38
Blue 1342244188_L2_5 120 200 35000 Training 4800 40
Blue 1342251926_L2_5 115 200 30000 Training 4370 38
Blue 1342239278_L2_5 115 100 30000 Training 4830 42
Blue 1342255011_L2_5 115 500 40000 Training 4255 37
Blue 1342244168_L2_5 100 2000 40000 Training 2800 28
Blue 1342255006_L2_5 90 800 40000 Training 3060 34
Blue 1342189109_L2_5 90 500 40000 Training 3330 37
Blue 1342244166_L2_5 80 600 40000 Training 2880 36
Blue 1342204366_L2_5 80 800 40000 Training 2720 34
Blue 1342255009_L2_5 75 1000 40000 Training 2400 32
Blue 1342214761_L2_5 75 600 40000 Training 2700 36
Blue 1342204368_L2_5 25 400 40000 Training 950 38
Blue 1342244831_L2_5 125 200 40000 Validation 5250 42
Blue 1342205078_L2_5 120 200 30000 Validation 4560 38
Blue 1342188088_L2_5 90 60 15000 Validation 3330 37
Blue 1342214510_L2_5 55 400 40000 Validation 2090 38
Blue 1342185644_L2_5 60 80 14000 Test 2100 35
Blue 1342257384_L2_5 85 250 40000 Test 3485 41
Blue 1342257382_L2_5 115 1000 40000 Test 3680 32
Blue 1342244170_L2_5 100 600 40000 Test 3600 36
Blue 1342186121_L2_5 75 350 20000 Test 2325 31
Blue 1342184484_L2_5 80 600 40000 Test 2880 36
Green 1342189661_L2_5 210 150 25000 Training 7980 38
Green 1342205047_L2_5 180 200 30000 Training 6840 38
Green 1342187331_L2_5 170 200 40000 Training 7140 42
Green 1342189845_L2_5 150 200 30000 Training 5700 38
Green 1342188096_L2_5 140 150 25000 Training 5320 38
Green 1342223213_L3 60 40 12000 Training 2100 35
Green 1342187170_L2_5 115 250 25000 Validation 4025 35
Green 1342190275_L2_5 70 80 18000 Validation 2660 38
Green 1342198592_L2_5 70 100 25000 Validation 2800 40
Green 1342198246_L2_5 115 150 30000 Test 4600 40
Green 1342197280_L2_5 75 150 20000 Test 2700 36
Green 1342190273_L2_5 70 80 25000 Test 2870 41
Green 1342239263_L2_5 65 100 20000 Test 2470 38
Red 1342257382_L2_5 90 600 40000 Test 3240 36
Red 1342188088_L2_5 80 300 15000 Test 2320 29
Red 1342257384_L2_5 75 300 40000 Test 3000 40
Red 1342186121_L2_5 70 400 20000 Test 2100 30
Red 1342184484_L2_5 65 600 40000 Test 2340 36
Red 1342185644_L2_5 55 125 14000 Test 1815 33
 [^10] The columns from left to right represent: the band, observation ID (field), number of sources per simulation injected into each simulation of this field, minimum and maximum flux density levels, dataset split (Training, Validation, or Test), total number of sources (calculated as the number of simulations multiplied by the number of sources per simulation), and finally, the number of simulations, which corresponds to the range of flux levels between the minimum and maximum flux densities.